OpenStudy (anonymous):

Points E(8, 14) and F(3, 2) on a coordinate grid represent side EF of triangle EFG. What is the length of side EF of the triangle? Answer 4 units 6 units 13 units 18 units

OpenStudy (anonymous):

13 units

hero (hero):

\[a^2 + b^2 = c^2 \]\[5^2 + 12^2 = c^2\]\[25 + 144 = c^2\]\[169 = c^2\]\[\sqrt{169} = c\]\[13 = c\]